# Leverage Knowledge Graph and Large Language Model for Law Article Recommendation: A Case Study of Chinese Criminal Law

### v1

@@ 1 Introduction
In a modern rule-of-law society, the efficiency of court judgments is crucial for ensuring economy growth and social stability [39, 53, 41]. Meanwhile, [60] revealed that the current backlog of cases in courts is enormous, which means the workload of judicial personnel is heavy. [59] showed that there is a great need for intelligent tools to assist judicial personnel in improving the efficiency of case adjudication, since intelligent tools boost judicial efficiency by categorizing similar cases, speeding up research, and staying updated with new laws, ensuring faster and more consistent decisions.
@@
Some researchers utilized Large Language Model (LLM) to complete the task of law article recommendation [61, 50, 56]. When engaging with extensive LLMs, users are enabled to not only retrieve the numeral and textual format of law article but also to investigate the rationale behind the applicability of these statutes to the specific case by posing supplemental inquiries. Moreover, users have the capability to probe into the rationale for the non-selection of other analogous provisions. However, [37] found that directly using LLMs is prone to hallucinations, generating incorrect law article information. Therefore, it is necessary to introduce a fully accurate legal information and case knowledge base for auxiliary analysis.
@@
To address the above issues, we introduce knowledge graph technology and large language model technology to improve the output accuracy in law article recommendation tasks. Initially, we pre-design the schema of a Case-Enhanced Law Article Knowledge Graph (CLAKG) capable of simultaneously storing current law articles and historical cases information as a database. On this basis, we also introduce an automated CLAKG construction method based on LLM. Furthermore, we design a closed-loop human-machine collaboration method to complete the law article recommendation task. To demonstrate the effectiveness of our proposed method, we constructed a dataset of Chinese criminal law judgments. We compared the method introduced in this paper with approaches based on BERT, methods utilizing LLM, and the LLM approach based on TFIDF.
@@
The remainder of this paper is organized as follows. Section 2 proposes a closed-loop law article recommendation framework. Section 3 utilizes the method from the second section to construct the CLAKG and compares the accuracy of this method with other baseline approaches. Lastly, the conclusions and future work are highlighted in Section 4.
@@ 2.1 Case-Enhanced Law Article Knowledge Graph Construction
(a) Schema of CLAKG. Ellipses of different colors represent
@@
(b) Example of CLAKG.
Figure 2: Overall schema and instantiation example.
@@
As shown in Figure 2(a), the Case-Enhanced Law Article Knowledge Graph (CLAKG) is composed of the Law Article Knowledge Graph (LAKG) and the Adjudicated Case Knowledge Graph (ACKG). The LAKG includes original article of law, law article id, and key information. The key information is extracted from the original article of law using a LLM. For a more detailed instruction of LLM prompts in the task of retrieving law article key information or a specific example, we direct the reader to Section 1.1 of supplementary material. The ACKG contains the name of a case, time of the court session, prosecution reason, and the specifics of a case. The description of the case process is derived from the content of the judgment documents processed using a LLM. For a more detailed instruction of LLM prompts in the task of retrieving case process, we direct the reader to Section 1.2 of supplementary material.
Based on the reference law articles written in the judgment documents, it is possible to link the name of a case node with law article id nodes. We analyze the correlation between case information and the corresponding key information of applicable law articles using a LLM and ultimately select up to five of the most relevant keywords for connection. For a more detailed instruction of LLM prompts in the task of matching case information to the corresponding key information of applicable law articles, we direct the reader to Section 1.3.2 of supplementary material. To ensure the consistency and accuracy of the knowledge graph. We ignore any keywords that do not exist in the graph output by the large language model. Tables 1 and 2 provide the details of all types of nodes and relationships in the CLAKG.
@@ 2.2.1 Graph Embedding on CLAKG with RGCN
1:Initialize: law_article_distances = [] , law_article_set = {} # Initializes a list to store distances and corresponding law articles and a set to store candidate law articles
@@
3:  Case_set.add_node($x\mid x-[\text{key}]\rightarrow e$) # $x$ is ‘Original article of penal law’ node
@@
6:  $\text{distance}=0$ # Initializes the distance score
@@
8:   $\text{distance}\mathrel{+}=\operatorname{dis}(\text{emb}[\text{law\_article}],\text{emb}[\text{e}])$ # Calculate the distance score
9:    where $\operatorname{dis}(a,b)=\frac{a\cdot b}{\|a\|\times\|b\|}$ # Cosine similarity formula
@@
14:return law_article_distances[:5] # Output the top five candidate law articles
Figure 4: Flowchart of RGCN training process.
$f(s,r,o)=e_{s}^{T}R_{r}e_{o}$ (2)
$\operatorname{Loss}=$ $-\frac{1}{(1+\omega)|\hat{\varepsilon}|}\sum_{(s,r,o,y)\in\mathcal{T}}\Bigg[y\log l(f(s,r,o))$ (3)
$+(1-y)\log\left(1-l(f(s,r,o))\right)\Bigg]$
$l(x)=\frac{1}{1+e^{-x}}=\frac{e^{x}}{e^{x}+1}$ (4)
2.2.2 Matching Key Information Nodes in CLAKG with LLM
In this section, we aim to match the relevant key information nodes within CLAKG based on current case information provided by users. To facilitate this process, we have designed prompts for the LLM, which include the expert role, task description, details of the newly reported case, known key information nodes, and example outputs. For a more detailed instruction of LLM prompts in the task of matching key information nodes in the CLAKG related to new case information, we direct the reader to Section 2.1 of supplementary material.
2.2.3 Law Article and Case Retrieval Algorithm
@@
The proposed method for retrieving candidate law article involves traversing all ‘key information’ nodes obtained in 2.2.2 and retrieving the ‘Original article of law’ nodes that are connected to the ‘key information’ nodes via ‘key’ relationships. These nodes are then aggregated into a set container, with all ‘Original article of law’ nodes in the set representing the candidate law article. Given that the number of candidate law articles in the set container may be substantial, it is essential to select those law articles most analogous to the newly reported case for further analysis. To this end, each node in the container is evaluated by calculating the distance scores (cosine similarity) between the ‘Original article of penal law nodes’ embedding vector and the multiple ‘key information’ nodes embedding vector obtained in 2.2.1.
@@ 2.2.4 LLM-Enhanced Law Article Recommendation
For a more detailed instruction of LLM prompts in the task of law article recommendation, we direct the reader to Section 2.3 of supplementary material. After the user describes the case information to the system, the system returns the results generated by the LLM. The user can further inquire about the results provided by the LLM.
@@ 3.1 Datasets
The criminal case data analyzed in this study were obtained from the China Judgments Online [58] platform. Figure 5 presents one of the cases. In light of the fact that the Amendment (XI) to the Criminal Law of the People’s Republic of China came into force on March 1, 2021, all judicial decisions used herein pertain to cases adjudicated after this date. During this period, due to concerns over information security and privacy protection, the number of publicly available criminal case judgments was limited. As a result, certain law articles corresponded with only a limited number of cases, while some provisions had no corresponding cases at all. We curated a dataset comprising 1,170 case records, which collectively address law articles from 85 articles within the Specific Provisions. To optimize the dataset, we divided it in a manner that maximized the diversity of law articles represented in the test set. Specifically, 997 case records were designated for the training set and 173 for the test set. The training set encompassed 49 distinct law articles, whereas the test set covered 85 law articles (zoro_shot_num = 36). The Adjudicated Case Knowledge Graph was subsequently constructed based on the training data, following the approach outlined in Section 2.1.
@@ 3.2 CLAKG Construction and Graph Embedding
We employed the method described in Section 2.1 to integrate the LAKG and the ACKG, resulting in the CLAKG. The types and numbers of nodes in CLAKG are presented in Table 3, while the types and numbers of relationships are detailed 4.
Table 3: Entity types and numbers in CLAKG.
Entity Type Number
Original article of law 452
Key information 3405
Law article id 452
The name of a case 997
Time of the court session 997
Prosecution reason 997
The specifics of the case 997
Table 4: Relation types and numbers in CLAKG.
Relation Type Number
Key 7854
Id 452
Agree with 997
Applicable law 4118
Occur in time 997
Reason 997
Detail 997
@@ 3.3 Case Study
We illustrate the law recommendation process based on LLM and CLAKG using the case of ”Zhang Yue’s offenses of bribery and abuse of power” as an example. Initially, the user inputs the details of a new case:
@@
The system then retrieves keywords pertinent to the case using the methodology outlined in Section 2.2.2, identifying the following terms: [‘accepting bribes’, ‘abuse of power’, ‘bribery’]. Subsequently, the system locates the corresponding node IDs within the CLAKG and obtains the graph embedding vectors for these nodes as described in Section 3.2. Utilizing the historical case information and the candidate law articles retrieval algorithm specified in Section 2.2.3, the system identifies the candidate law articles and their associated cases, as depicted in Section 2.2 of supplementary material.
The system consolidates the above information into a prompt using the method described in Section 2.2.4, and submits it to the LLM to carry out the law recommendation task. Ultimately, the LLM outputs “Article 385” as the recommended law article. The result produced by the LLM is consistent with the outcome recorded in the court judgment.

### v2

@@ Abstract
keywords:
law article recommendation , Large Language Model , Knowledge Graph
[…]
[label1]organization=Global Institute of Future Technology, Shanghai Jiao Tong University, city=Shanghai, postcode=200240, state=Shanghai, country=China
[…]
[label2]organization=Donghua University, city=Shanghai, postcode=201620, state=Shanghai, country=China
{highlights}
Proposed a novel Case-Enhanced Law Article Knowledge Graph (CLAKG) that integrates law articles and historical case data, enhancing the accuracy of law article recommendations.
Introduced an automated CLAKG construction method that uses a Large Language Model (LLM) to reduce manual input and improve scalability. Developed a closed-loop recommendation system, leveraging LLMs and CLAKG for more accurate law article recommendations while mitigating common issues like hallucinations in LLM outputs.
Achieved significant improvement in accuracy: The proposed method improved law article recommendation accuracy from 0.549 to 0.694, outperforming baseline models such as BERT, DPCNN, TFIDF-RAG, Graph-RAG and Light-RAG.
@@ 1 Introduction
In a modern rule-of-law society, the efficiency of court judgments is crucial for ensuring economic growth and social stability[1, 2, 3]. Meanwhile, [4] revealed that the current backlog of cases in courts is enormous, which means the workload of judicial personnel is heavy. [5] showed that there is a great need for intelligent tools to assist judicial personnel in improving the efficiency of case adjudication, since intelligent tools boost judicial efficiency by categorizing similar cases, speeding up research, and staying updated with new laws, ensuring faster and more consistent decisions.
@@
Some researchers utilized Large Language Models (LLMs) to complete the task of law article recommendation[12, 13, 14]. By using a large LLM, users can not only retrieve the law article in numeric and textual form, but also investigate the rationale behind its applicability to the case by asking follow-up questions. Moreover, users have the capability to probe into the rationale for the non-selection of other analogous provisions. However, [15] found that directly using LLMs is prone to hallucinations, generating incorrect law article information. Therefore, it is necessary to introduce a fully accurate legal information and case knowledge base for auxiliary analysis.
@@
To address the above issues, we introduce knowledge graph technology and large language model technology to improve the output accuracy in law article recommendation tasks. Initially, we pre-design the schema of a Case-Enhanced Law Article Knowledge Graph (CLAKG) capable of simultaneously storing current law articles and historical cases information as a database. On this basis, we also introduce an automated CLAKG construction method based on LLM. Furthermore, we design a closed-loop human-machine collaboration method to complete the law article recommendation task. To demonstrate the effectiveness of our proposed method, we constructed a dataset of Chinese criminal law judgments. We compared the method introduced in this paper with approaches based on BERT[23], GRU[8], DPCNN[11], and several RAG approaches such as TFIDF-RAG[21], Graph-RAG[27] and Light-RAG[28]. Finally, we conducted an ablation study of our method.
@@
The remainder of this paper is organized as follows. Section 2 reviews related works about law article recommendation, RAG technology and knowledge graphs in legal domain. Section 3 proposes a closed-loop law article recommendation framework. Section 4 utilizes the method from the third section to construct the CLAKG and compares the accuracy of this method with other baseline approaches. Lastly, the conclusions and future work are highlighted in Section 5.
2 Related Work
2.1 Law Article Recommendation
Early approaches to law article recommendation treated the task as a text classification or retrieval problem. Classical methods relied on keyword matching or manual feature engineering, but these struggled with the complex language of statutes. With the rise of deep learning, researchers applied neural models like LSTMs[6] and CNNs[7] to match case fact descriptions with relevant law articles. These models improved over bag-of-words baselines by capturing sequential and local context features. Transformer-based language models further advanced the field of BERT[29] and its legal-domain adaptations (e.g., LegalBERT[30]) enabled richer semantic representations of legal text. For example, Chalkidis et al.[31] showed that a BERT-based classifier can predict applicable European human rights articles more accurately than traditional models, while Aletras et al.[32] demonstrated early on the feasibility of data-driven prediction of court decisions and their pertinent articles. Despite these gains, most prior models formulate law article recommendation as mapping facts to article IDs, without directly leveraging the content of the law articles themselves. This focus solely on fact-to-ID correspondence makes them heavily data-dependent and sensitive to label imbalance in training sets[33]. In practice, it also limits interpretability – the models do not explain why a recommended article is relevant, since the law articles are not explicitly used in the reasoning process. These limitations have motivated explorations into techniques that can incorporate external legal knowledge and provide better justification for recommendations.
2.2 LLMs and Retrieval-Augmented Generation Techniques
The advent of large language models (LLMs), such as GPT-style transformers, has opened new avenues for law article recommendation. Rather than training a specific classifier, one can prompt an LLM with a case’s facts and ask for relevant law articles. Recent works have started to investigate this approach[12, 13, 14], noting that LLMs can, in principle, leverage their vast pretrained knowledge to identify relevant law articles and even generate explanations. In particular, Shui et al.[14] evaluate various state-of-the-art LLMs on legal judgment prediction subtasks (including relevant law article prediction), finding that while LLMs exhibit promising zero-shot capabilities, they still lag behind task-specific models on accuracy. A major concern with directly using LLMs for legal guidance is hallucination: the model may produce plausible-sounding but incorrect or nonexistent legal references. Dahl et al.[15] provide a sobering analysis of this issue, showing that even advanced LLMs frequently fabricate legal citations or statutes, which is unacceptable in high-stakes domains like law. To mitigate such hallucinations, researchers are turning to Retrieval-Augmented Generation (RAG) techniques. RAG combines LLMs with information retrieval: before or during text generation, the system retrieves relevant documents (e.g. law articles or case precedents) and conditions the LLM on that retrieved evidence. By grounding responses in actual law articles, RAG can significantly improve factual accuracy and trustworthiness of the model’s output. This approach was originally popularized in open-domain QA systems[16] and has since been applied to legal NLP as well. For instance, a RAG-based legal assistant might fetch the text of candidate law articles given a case description, and the LLM then quotes or summarizes those law articles when recommending applicable law, thereby rooting its answer in verifiable sources. Key components of RAG include a retriever and a generator. In the legal domain, classic IR methods like TF–IDF[21] or BM25[22] have been used to retrieve statute candidates by keyword overlap, while more recent approaches use neural retrievers (e.g. dense embeddings). Borgeaud et al.[23] showed that even a frozen BERT-based retriever can substantially improve generation quality by supplying relevant text from a large corpus. However, naive retrieval based on surface text may miss conceptual matches. Legal phrasing varies, and a model might need to retrieve statutes using synonymous terms or implied connections, something neural retrieval aims to handle better than lexical methods. Overall, RAG represent a promising direction for law article recommendation, as they combine the reasoning ability of generative models with the grounding of information retrieval, thereby reducing hallucinations and improving the transparency of the recommendation.
2.3 Knowledge Graphs in the Legal Domain
Beyond unstructured text retrieval, Knowledge Graphs (KGs) have emerged as valuable resources for legal information retrieval and recommendation. A legal knowledge graph represents entities such as law articles, cases, legal concepts, and their relationships (e.g., a case cites a statute, or a chapter includes certain articles) in a structured form. This graph structure can complement text-based methods by capturing the domain’s ontology and the interconnections between law articles. Prior works have constructed legal knowledge graphs to assist various tasks. For example, law-specific knowledge bases have been used to integrate charge definitions and article relationships into prediction models, yielding performance gains and better interpretability in tasks like judgment prediction[34, 35]. More recently, Chaudhri et al.[24] argued that KGs are well-suited as the underlying database for RAG systems: instead of retrieving free text, one can query a knowledge graph to retrieve facts or linked nodes, ensuring that the generation is grounded in a consistent knowledge structure. In the legal domain, a KG could be queried to find statutes related to certain legal concepts or past cases with similar fact patterns. One advantage of KGs is that they store high-level associations beyond lexical similarity – for instance, a graph can link a section of law to broader categories or to precedent cases, enabling the recommendation system to recall relevant law articles even if a case’s description uses different wording. There is growing interest in integrating KGs with LLMs[25]. Pan et al. [25] outline a roadmap for unifying the strengths of symbolic knowledge graphs and distributed language models, noting that KGs can provide factual grounding for LLMs while LLMs can help populate and expand KGs from raw texts. In the context of law, researchers have begun designing hybrid systems that use KGs alongside LLMs. Chen et al.[33], for instance, construct a case-enhanced law article KG that links statutes with relevant past cases, and then use an LLM-based pipeline to recommend articles by traversing the graph for pertinent nodes (law articles) and using the LLM to generate the final recommendations and explanations. Such approaches have demonstrated improved accuracy over text-only baselines, highlighting that the structured knowledge in KGs can effectively guide the retrieval and generation process. Meanwhile, Edge et al.[26] explore automatic construction of knowledge graphs using LLMs (generating graph triples from unstructured legal text) and using those graphs to enhance answer generation. This suggests a future where legal KGs and LLMs continuously enrich each other: the KG offers a source of truth to reduce LLM errors, and the LLM helps keep the KG up-to-date with new insights. The integration of knowledge graphs in legal AI is still an emerging area, and challenges remain, such as ensuring the completeness and currency of the graph. Nonetheless, the combination of KGs with retrieval-augmented LLMs holds great promise for building more reliable and interpretable law article recommendation systems in the coming years.
@@ 3.1 Case-Enhanced Law Article Knowledge Graph Construction
Figure 2: Schema of CLAKG. Ellipses of different colors represent
@@
Figure 3: Example of CLAKG.
@@
As shown in Figure 2, the Case-Enhanced Law Article Knowledge Graph (CLAKG) is composed of the Law Article Knowledge Graph (LAKG) and the Adjudicated Case Knowledge Graph (ACKG). The LAKG includes original article of law, law article id, and key information. The key information is extracted from the original article of law using a LLM. The ACKG contains the name of a case, time of the court session, prosecution reason, and the specifics of a case. The description of the case process is derived from the content of the judgment documents processed using a LLM.
Based on the reference law articles written in the judgment documents, it is possible to link the name of a case node with law article id nodes. We analyze the correlation between case information and the corresponding key information of applicable law articles using a LLM and ultimately select up to five of the most relevant keywords for connection. To ensure the consistency and accuracy of the knowledge graph, we ignore any keywords that do not exist in the graph output by the large language model. Tables 1 and 2 provide the details of all types of nodes and relationships in the CLAKG.
@@ 3.2.1 Graph Embedding on CLAKG with RGCN
Figure 5: Flowchart of RGCN training process.
$f(s,r,o)=e_{s}^{T}R_{r}e_{o}$ (2)
$\operatorname{Loss}=$ $-\frac{1}{(1+\omega)|\hat{\varepsilon}|}\sum_{(s,r,o,y)\in\mathcal{T}}\Bigg[y\log l(f(s,r,o))$ $+(1-y)\log\left(1-l(f(s,r,o))\right)\Bigg]$ (3)
$l(x)=\frac{1}{1+e^{-x}}=\frac{e^{x}}{e^{x}+1}$ (4)
3.2.2 Matching Key Information Nodes in CLAKG with LLM
In this section, we aim to match the relevant key information nodes within CLAKG based on current case information provided by users. To facilitate this process, we have designed prompts for the LLM, which include the expert role, task description, details of the newly reported case, known key information nodes, and example outputs. The prompts for LLM in the task of matching key information nodes in the CLAKG related to new case information is shown in Table 3.
Table 3: Prompts for LLM in the task of matching key information nodes in the CLAKG related to new case information.
Component Details
Expert Role Expert in law article analysis
Task Description Given the available new case information and the following key information nodes, please output 080-8 key information nodes most relevant to the case. If you believe there are fewer than 8 relevant key information nodes, there is no need to force the list to reach 8 nodes.
New Case Information Description of the new case input by the user
Key Information Nodes all key information nodes in the CLAKG
Precautions Please directly output the key information nodes, separated by semicolons (;) without including any additional content (including unnecessary punctuation marks).
Output Example Public and private property; multiple thefts; sale; committing a crime; causing damage; seriously disrupting public order; multiple thefts; lawfully performing duties
3.2.3 Law Article and Case Retrieval Algorithm
@@
1:Initialize: law_article_distances = [] , law_article_set = {} $\triangleright$ Initializes a list to store distances and corresponding law articles and a set to store candidate law articles
@@
3:  Case_set.add_node($x\mid x-[\text{key}]\rightarrow e$) $\triangleright$ $x$ is ‘Original article of penal law’ node
@@
6:  $\text{distance}=0$ $\triangleright$ Initializes the distance score
@@
8:   $\text{distance}\mathrel{+}=\operatorname{dis}(\text{emb}[\text{law\_article}],\text{emb}[\text{e}])$ $\triangleright$ Calculate the distance score
9:    where $\operatorname{dis}(a,b)=\frac{a\cdot b}{\|a\|\times\|b\|}$ $\triangleright$ Cosine similarity formula
@@
14:return law_article_distances[:5] $\triangleright$ Output the top five candidate law articles
@@
The proposed method for retrieving candidate law article involves traversing all ‘key information’ nodes obtained in 3.2.2 and retrieving the ‘Original article of law’ nodes that are connected to the ‘key information’ nodes via ‘key’ relationships. These nodes are then aggregated into a set container, with all ‘Original article of law’ nodes in the set representing the candidate law article. Given that the number of candidate law articles in the set container may be substantial, it is essential to select those law articles most analogous to the newly reported case for further analysis. To this end, each node in the container is evaluated by calculating the distance scores (cosine similarity) between the ‘Original article of penal law nodes’ embedding vector and the multiple ‘key information’ nodes embedding vector obtained in 3.2.1.
@@ 3.2.4 LLM-Enhanced Law Article Recommendation
After the user describes the case information to the system, the system returns the results generated by the LLM. The user can further inquire about the results provided by the LLM.
@@ 4.1 Datasets
The criminal case data analyzed in this study were obtained from the China Judgments Online [39] platform. Figure 6 presents one of the cases. In light of the fact that the Amendment (XI) to the Criminal Law of the People’s Republic of China came into force on March 1, 2021, all judicial decisions used herein pertain to cases adjudicated after this date. During this period, due to concerns over information security and privacy protection, the number of publicly available criminal case judgments was limited. As a result, certain law articles corresponded with only a limited number of cases, while some provisions had no corresponding cases at all. We curated a dataset comprising 1,170 case records, which collectively address law articles from 85 articles within the Specific Provisions. To optimize the dataset, we divided it in a manner that maximized the diversity of law articles represented in the test set. Specifically, 997 case records were designated for the training set and 173 for the test set. The training set encompassed 49 distinct law articles, whereas the test set covered 85 law articles (zoro shot number = 36). The Adjudicated Case Knowledge Graph was subsequently constructed based on the training data, following the approach outlined in Section 3.1.
@@ 4.2 CLAKG Construction and Graph Embedding
We employed the method described in Section 3.1 to integrate the LAKG and the ACKG, resulting in the CLAKG. The types and numbers of nodes as well as relationships in CLAKG are presented in Table 4.
Table 4: Entity and relation types with their Numbers in CLAKG.
Entity Type Number Relation Type Number
Original article of law 452 Key 7854
Key information 3405 Id 452
Law article id 452 Agree with 997
The name of a case 997 Applicable law 4118
Time of the court session 997 Occur in time 997
Prosecution reason 997 Reason 997
The specifics of the case 997 Detail 997
Figure 7: Train loss and AUC over epochs.
@@ 4.3 Case Study
We illustrate the law article recommendation process based on LLM and CLAKG using the case of ”Zhang Yue’s offenses of bribery and abuse of power” as an example. Initially, the user inputs the details of a new case:
@@
The system then retrieves keywords pertinent to the case using the methodology outlined in Section 3.2.2, identifying the following terms: [‘accepting bribes’, ‘abuse of power’, ‘bribery’]. Subsequently, the system locates the corresponding node IDs within the CLAKG and obtains the graph embedding vectors for these nodes as described in Section 4.2. Utilizing the historical case information and the candidate law articles retrieval algorithm specified in Section 3.2.3, the system identifies the candidate law articles and their associated cases.
The system consolidates the above information into a prompt using the method described in Section 3.2.4, and submits it to the LLM to carry out the law article recommendation task. Ultimately, the LLM outputs “Article 385” as the recommended law article. The result produced by the LLM is consistent with the outcome recorded in the court judgment.